OpenStudy (anonymous):

2^2X+1=2^X-3 solve the exponential equation

OpenStudy (anonymous):

take the ln or log of both sides. then the exponent can be brought out in front. Also ln 1 = 0 and log 1 = 0

OpenStudy (anonymous):

uh...lol what

OpenStudy (anonymous):

\[\large 2^{2x}+1=2^{x-3}\]\[\large \log 2^{2x}+\log 1=\log 2^{x-3}\]\[\large 2x \log 2+\log 1=(x-3) \log 2\]\[\log_{10}1=x\]\[10^x=1\]\[x=0 \rightarrow \log 1=0\]\[\large 2x \log 2=(x-3) \log 2\]You can take it from here